OpenStudy (anonymous):

can u check if i am right? differentiate f(x)=x^2e^-3x

OpenStudy (anonymous):

f'(x)=2xe^-3x-3x^2e^-3x

OpenStudy (anonymous):

more clear

OpenStudy (anonymous):

f'(x)=(2x e^-3x)-(3x^2 e^-3x)

OpenStudy (anonymous):

\[f(x)=x^2e^{-3x}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

product rule problem. \[(fg)'=f'g+g'f\] use \[f(x)=x^2,f'(x)-2x,g(x)=e^{-3x},g'(x)=-3e^{-3x}\]

OpenStudy (anonymous):

so im right

OpenStudy (anonymous):

gives \[2xe^{-3x}-3x^2e^{-3x}\]

OpenStudy (anonymous):

yes you are right.

OpenStudy (anonymous):

OHHHHH i thought it was \[f(x)=x ^{2e ^{-3x}}\] lololol

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

jiji

OpenStudy (anonymous):

thats some majorrrr chain ruleee bro